# Not All Contexts Are Equal: Teaching LLMs Credibility-aware Generation

### v1

@@ Abstract
The rapid development of large language models has led to the widespread adoption of Retrieval-Augmented Generation (RAG), which integrates external knowledge to alleviate knowledge bottlenecks and mitigate hallucinations. However, the existing RAG paradigm inevitably suffers from the impact of flawed information introduced during the retrieval phrase, thereby diminishing the reliability and correctness of the generated outcomes. In this paper, we propose Credibility-aware Generation (CAG), a universally applicable framework designed to mitigate the impact of flawed information in RAG. At its core, CAG aims to equip models with the ability to discern and process information based on its credibility. To this end, we propose an innovative data transformation framework that generates data based on credibility, thereby effectively endowing models with the capability of CAG. Furthermore, to accurately evaluate the models’ capabilities of CAG, we construct a comprehensive benchmark covering three critical real-world scenarios. Experimental results demonstrate that our model can effectively understand and utilize credibility for generation, significantly outperform other models with retrieval augmentation, and exhibit resilience against the disruption caused by noisy documents, thereby maintaining robust performance. Moreover, our model supports customized credibility, offering a wide range of potential applications.  [^1] Our code, benchmark, and models are available at https://github.com/panruotong/CAG
@@ 1 Introduction
However, RAG for large language models remains significantly impacted by flawed information. This is mainly because the retrieval process often provides noisy, outdated, and incorrect contexts which adversely affects RAG, substantially reducing its effectiveness. Specifically, previous research Shi et al. (2023a); Chen et al. (2023a) has found that LLMs are highly sensitive to noise, which impacts LLMs’ capacity to discern and trust accurate information, ultimately affecting the outcomes they generate. Furthermore, due to the temporal insensitivity of LLMs  (Su et al., 2022; Zhao et al., 2024), these models struggle to discern outdated information solely based on their internal knowledge. More critically, because LLMs are trained on extensive collections of historical text, there’s an inherent risk that outdated information will align with the models’ internal knowledge bases. This alignment can inadvertently encourage LLMs to favor and perpetuate outdated information. Besides, the prevalence of misinformation on the current web poses a significant challenge for large models, which struggle to identify misinformation using only their inherent knowledge (Xie et al., 2023; Pan et al., 2023). This difficulty makes them susceptible to misinformation, leading to the generation of incorrect answers. Therefore, flawed information, characterized by noisy, outdated, and incorrect information, has substantial negative effects on RAG.
@@
From a cognition perspective, a common approach humans adopt to combat flawed information is to assess the credibility of external information Burgoon et al. (2000). Specifically, information that is current, evaluated, and sourced from highly credible origins is typically regarded as more timely, accurate, and reliable. Motivated by this, we introduce Credibility-aware Generation (CAG), a universally applicable framework designed to address flawed information encountered during RAG. At its core, CAG seeks to equip models with the ability to discern and process information based on credibility. By attributing varying levels of credibility to information differing in relevance, temporal context, and source, and integrating these credibility indicators into the generative processes of LLMs, CAG significantly mitigates the issues arising from flawed information.
@@
Unfortunately, we have discovered that existing LLMs are not inherently sensitive to directly provided credibility information in the prompt. This deficiency restricts their capacity to optimally employ credibility for discerning and processing information. To endow models with the capability of CAG, we propose a novel data transformation framework based on existing Question Answering (QA) and dialogue datasets. This framework transforms the data into a format that incorporates credibility and can be utilized to guide model credibility-based generation, thereby training the model to utilize credibility in addressing flawed information. Specifically, our data transformation process comprises two core steps: 1) Multi-granularity credibility annotation, which assigns credibility to text units at both document and sentence levels by dividing retrieved documents into varying granularities. 2) Credibility-guided explanation generation, which prompts LLMs to generate credibility-guided explanations given questions, retrieved documents with credibility annotation and golden answers. Finally, we utilize instruction fine-tuning to train the model, enabling it to generate responses based on credibility.
@@
To rigorously assess the ability of the model’s credibility-aware generation in managing flawed information, we have developed a comprehensive benchmark encompassing various real-world scenarios, including open-domain QA, time-sensitive QA, and misinformation-polluted QA. In these settings, several indicators, including retrieval relevance, temporal validity, and source authority, are considered as the given credibility measurement. The main goal of this benchmark is to measure how well a model can answer questions when given the context documents with corresponding credibility annotations. Experimental results on multiple datasets across multiple scenarios demonstrate the efficacy of our approach in leveraging credibility information. Our model significantly outperforms various prevalent RAG approaches applied to both open and closed-source LLMs of diverse scales. Additionally, it exhibits robust resilience against noisy documents, maintaining high performance levels even as alternative strategies experience sharp deteriorations. All these results verify the effectiveness of the proposed CAG framework and corresponding training algorithm.
@@
Figure 2: Overview of data transformation framework. The training data is constructed by first assigning credibility to contexts via multi-granularity credibility annotation (§3.1) and prompting LLM to produce credibility-guided explanations (§3.2). The processed data is used to instruction fine-tuning (§3.3) to endow the model with the ability for Credibility-aware Generation.
The main contributions of this study are summarized as follows:
@@
  • We propose a novel data transformation framework that leverages existing QA and dialogue datasets. This framework transforms existing datasets into data annotated with credibility and guides models to generate responses based on credibility, thereby equipping the model with Credibility-aware Generation capability.
@@
  • Experimental evidences demonstrate that our model effectively understands and employs credibility to generate responses, significantly surpasses other RAG-based strategies, and maintain robustness despite the increasing noise in the context.
@@ 3.1 Multi-granularity Credibility Annotation
To cater to the varied requirements for credibility across different scenarios and enhance the model’s comprehension of credibility, we collect training data encompasses Open-domain QA, Machine Reading Comprehension (MRC) datasets, and dialogue datasets and propose a multi-granularity credibility annotation method.
First, we divide the retrieved documents to create a multi-granularity corpus, encompassing sentence and document levels. Then, the retriever assesses the match between each retrieval unit and the query, assigning a relevance score, and classifies documents into three levels: high, medium, and low, employing either equi-frequency or equi-distance segmentation. This approach of using levels instead of scores aims to simplify representation, thereby improving the model’s understanding and providing a certain degree of fault tolerance. Consequently, we gather about 15k training datasets, within which the contexts of the QA data are annotated with different granularities of credibility. The detailed composition of the training data is shown in the Appendix A.1.
@@ 3.2 Credibility-guided Explanation Generation
To facilitate the model’s comprehension and effective utilization of credibility, we employ LLM to generate explanations for the answers. These explanations stem from an analysis of both the content and credibility of the documents.
Given the limitations of current LLMs in comprehending credibility effectively, we utilize chain-of-thought prompts to guide LLM to generate credibility-guided explanations given questions, retrieved documents with credibility and golden answers. In this case, LLM is required to analyze document content and credibility, as well as the rationale for the derived answer after integrating all the information. Considering the accessibility and advanced capabilities of GPT-3.5, we elect to employ GPT-3.5 for the generation of explanations. In this way, we obtain high-quality, credibility-guided answer explanations. Then, we replace the original answers in the training data with credibility-guided explanations to form a novel QA dataset based on credibility. Within this dataset, the inputs include questions and external documents annotated with credibility, while the outputs are credibility-guided explanations.
@@ 3.3 Instruction Fine-tuning
Through the two steps above, the training dataset obtained contains credibility, which can be used to facilitate the model in gaining the capacity for CAG. We fine-tune the language model on this dataset to empower the model to discern and process information according to its credibility. As defined by Iyer et al. (2023), the loss function is as follows:
$\mathcal{L}(D_{\boldsymbol{x}};\theta)=-\sum_{i=1}^{N}\log p_{\theta}\left(y_{i}\mid\left[\boldsymbol{x},\left\{[c_{i},d_{i}]\right\}_{i=1}^{|D_{\boldsymbol{x}}|}\right],y_{<i}\right)$
@@ 4 Credibility-aware Generation Benchmark
  • Open-domain QA aims to accurately answer questions on a wide variety of topics without being limited to any particular area. It encompasses a broad spectrum of real-world applications that urgently require the integration of external knowledge to enhance the language model’s ability to address queries. This scenario thus necessitates the ability to effectively identify and process noise information.
  • Time-sensitive QA aims to give answers that are both correct and up-to-date, using the most recent information available. It poses a challenge for LLMs due to the rapidly changing nature of internet information. The inevitable inclusion of outdated documents when incorporating external sources further complicates matters. Even with timestamps provided for each document, LLMs might still erroneously prioritize outdated documents. This situation underscores the critical need for credibility in time-sensitive QA scenarios.
  • Misinformation polluted QA aims to tackle the issue of ensuring accurate answers in an environment polluted with misinformation. It presents a substantial challenge to LLMs, attributed to the misuse of LLMs and the consequent proliferation of fake news and misinformation (Zhuo et al., 2023; Pan et al., 2023). Consequently, in this scenario, it is crucial to take into account the quality and credibility of any introduced external information.
In the following, we will provide a detailed description of data construction for each scenario, and the statistics of CAGB are shown in the Table 1.
@@ 4.1 Open-domain QA
Our research utilizes data from several challenging QA datasets that have noise in the context they provide. HotpotQA (Yang et al., 2018b) and 2WikiMHQA (Ho et al., 2020) both require reasoning across multiple documents, and feature a high proportion of distracting documents (60%-80% and 80%, respectively). Importantly, the data we utilize from HotpotQA is extracted from the dev subset, whereas our training dataset is derived from the train subset. Musique (Trivedi et al., 2021) questions are of higher complexity, with up to 90% of distracting passages. ASQA (Stelmakh et al., 2022) is a long format QA dataset focused on ambiguous questions. RGB (Chen et al., 2023b) is a specialized benchmark used for evaluating the capabilities of models in the RAG scenario, with noise robustness being one of its aspects.
We assign credibility to the documents provided in the dataset in terms of retrieval relevance. The retriever is used to assign relevance scores based on the similarity of each document to the query. We then divide the documents into three categories based on the scores at equal intervals, including high, medium, and low. Subsequently, we divide the documents into three credibility levels, specifically high, medium, and low, based on relevance scores distributed at equal intervals.
@@
Table 1: Statistics of our Credibility-aware Generation Benchmark, which includes 8 dataset derived from 3 scenarios.
@@ 4.2 Time-sensitive QA
In order to construct a diverse, high-quality, and up-to-date news dataset, we annotate 500 time-sensitive questions along with their corresponding dates. These questions originate from real-world scenarios, including news question answering data from RealTime QA (Kasai et al., 2022), TAQA (Zhao et al., 2024), and questions adapted from news reports.
To simulate the simultaneous occurrence of varied information on the Internet, we utilize Google search API to gather three relevant documents and four distracting documents for each question, the latter being either irrelevant or outdated. This approach to document selection is crafted to emulate the intricate and heterogeneous nature of real-world information landscapes. Each news includes its publication date, thereby aiding in the evaluation of its timeliness. For document credibility annotation, we assess credibility based on relevance and time difference between the document’s publication and the posed question. Documents that are highly relevant to the given query and possess strong timeliness are considered to have high credibility, while documents that are irrelevant or outdated are regarded as having low credibility. We ensure the accuracy of answers by manually annotating.
@@
4.3 Misinformation Polluted QA
To simulate the generation of fake news, we first generated a claim utilizing the LLM, based on a question and a randomly selected incorrect option. This process transforms the question and incorrect option into a deceptive statement. Subsequently, we choose GPT-3.5 and Qwen as the generators for fake news, guiding them to generate documents of varying styles based on the claim, including news style and Twitter style. The prompts used and examples of the generated content are detailed in the Appendix A.4. The fictitious news articles produced by LLMs, due to their authenticity being deliberately compromised, are classified as having low credibility. Conversely, news articles from reputable news websites are considered to possess high credibility. We set the ratio of fake news at 0.5, 0.67, and 0.75 to evaluate the robustness of model against misinformation under various levels of pollution.
By simulating the process of generating fake news and annotating credibility based on source, we obtain a misinformation polluted QA dataset in the news domain, named NewsPollutedQA.
@@ 4.3 Misinformation Polluted QA
ChatGPT 0.334 0.368 0.194 0.404 0.773 0.646 0.231
LLaMA-2-7B 0.280 0.312 0.160 0.268 0.753 0.446 0.179
Vicuna-7B 0.278 0.296 0.116 0.358 0.677 0.632 0.229
Mistral-7B-Instruct 0.288 0.270 0.106 0.300 0.713 0.596 0.204
LLaMA-2-13B 0.366 0.370 0.164 0.321 0.820 0.484 0.204
LLaMA-2-70B 0.418 0.390 0.256 0.316 0.823 0.492 0.430
vanilla IFT 0.324 0.245 0.270 0.157 0.650 0.566 0.329
@@
ChatGPT 0.396 0.394 0.216 0.388 0.790 0.676 0.427
LLaMA-2-7B 0.302 0.376 0.200 0.375 0.730 0.510 0.265
Vicuna-7B 0.355 0.306 0.164 0.494 0.757 0.686 0.275
Mistral-7B-Instruct 0.338 0.334 0.166 0.414 0.790 0.694 0.373
LLaMA-2-13B 0.370 0.372 0.180 0.390 0.823 0.552 0.308
LLaMA-2-70B 0.422 0.504 0.320 0.388 0.833 0.563 0.306
vanilla IFT 0.348 0.448 0.276 0.304 0.663 0.658 0.344
@@
ChatGPT 0.422 0.388 0.182 0.440 0.807 0.704 0.408
LLaMA-2-7B 0.376 0.176 0.140 0.394 0.713 0.470 0.213
Vicuna-7B 0.349 0.266 0.091 0.490 0.740 0.654 0.279
Mistral-7B-Instruct 0.274 0.268 0.102 0.463 0.797 0.686 0.315
LLaMA-2-13B 0.360 0.384 0.164 0.399 0.803 0.497 0.227
@@
vanilla IFT 0.372 0.334 0.204 0.305 0.663 0.564 0.383
CAG-7B (ours) 0.509 0.578 0.340 0.496 0.897 0.810 0.442
CAG-13B (ours) 0.514 0.604 0.408 0.510 0.917 0.816 0.483
CAG-mistral-7B (ours) 0.502 0.540 0.384 0.505 0.900 0.836 0.613
Table 2: Model performance in our CAGB benchmark. The best/second best scores in each dataset are bolded/underlined. Our models substantially outperforms previous strategies across all 3 scenarios in CAGB. The results shown for EvolvingTempQA and RGB are at noise_ratio setting of 0.8, while NewsPollutedQA is at noise_ratio setting of 0.75.

### v3

@@ Abstract
The rapid development of large language models has led to the widespread adoption of Retrieval-Augmented Generation (RAG), which integrates external knowledge to alleviate knowledge bottlenecks and mitigate hallucinations. However, the existing RAG paradigm inevitably suffers from the impact of flawed information introduced during the retrieval phrase, thereby diminishing the reliability and correctness of the generated outcomes. In this paper, we propose Credibility-aware Generation (CAG), a universally applicable framework designed to mitigate the impact of flawed information in RAG. At its core, CAG aims to equip models with the ability to discern and process information based on its credibility. To this end, we propose an innovative data transformation framework that generates data based on credibility, thereby effectively endowing models with the capability of CAG. Furthermore, to accurately evaluate the models’ capabilities of CAG, we construct a comprehensive benchmark covering three critical real-world scenarios. Experimental results demonstrate that our model can effectively understand and employ credibility for generation, significantly outperform other models with retrieval augmentation, and exhibit robustness despite the increasing noise in the context. 111 Our code, benchmark, and models are available at https://github.com/panruotong/CAG
@@ 1 Introduction
However, RAG for large language models remains significantly impacted by flawed information. This is mainly because the retrieval process often provides noisy, outdated, and incorrect contexts which adversely affects RAG, substantially reducing its effectiveness. Specifically, previous research Shi et al. (2023a); Chen et al. (2023) has found that LLMs are highly sensitive to noise, which impacts LLMs’ capacity to discern and trust accurate information, ultimately affecting the outcomes they generate. Furthermore, due to the temporal insensitivity of LLMs  (Su et al., 2022; Zhao et al., 2024), these models struggle to discern outdated information solely based on their internal knowledge. More critically, because LLMs are trained on extensive collections of historical text, there’s an inherent risk that outdated information will align with the models’ internal knowledge bases. This alignment can encourage LLMs to favor and perpetuate outdated information. Besides, the prevalence of misinformation on the current web poses a significant challenge for large models, which struggle to identify misinformation using only their inherent knowledge (Xie et al., 2023; Pan et al., 2023). This difficulty makes them susceptible to misinformation, leading to the generation of incorrect answers. Therefore, flawed information, characterized by noisy, outdated, and incorrect information, has substantial negative effects on RAG.
@@
From a cognition perspective, a common approach humans adopt to combat flawed information is to assess the credibility of external information Burgoon et al. (2000). For humans, credibility refers to the acceptability based on the quality of the information, its source, and subjective evaluation. However, LLMs relying solely on internal knowledge to assess information credibility are unstable and unreliable (Xie et al., 2023). Therefore, we aim to guide LLMs’ acceptance of information by utilizing external indicators of credibility. We introduce Credibility-aware Generation (CAG), a universally applicable framework designed to address flawed information encountered during RAG. At its core, CAG seeks to equip models with the ability to discern and process information based on credibility. By assigning different credibility to information based on its relevance, timeliness, and the reliability of its source, and explicitly distinguishing them in the input, CAG significantly mitigates the issues arising from flawed information.
@@
Unfortunately, we have discovered that existing LLMs are not inherently sensitive to directly provided credibility in the prompt. This deficiency restricts their capacity to optimally employ credibility for discerning and processing information. To endow models with the capability of CAG, we propose a novel data transformation framework. This framework transforms existing Question Answering (QA) datasets into data that integrates credibility, which can be employed to guide the model for credibility-based generation. Specifically, our process comprises two core steps: 1) Multi-granularity credibility annotation, which assigns credibility to text units at both document and sentence levels by dividing retrieved documents into varying granularities. 2) Credibility-guided explanation generation, which prompts LLMs to generate credibility-guided explanations given questions, retrieved documents with credibility annotation and golden answers. Finally, we employ instruction fine-tuning to train the model, enabling it to generate responses based on credibility.
@@
To rigorously assess the ability of the model’s credibility-aware generation in managing flawed information, we construct a comprehensive benchmark encompassing various real-world scenarios, including open-domain QA, time-sensitive QA, and misinformation polluted QA. In this benchmark, retrieval relevance, timeliness, and source authority are regarded as established measures of credibility. Experimental results on multiple datasets across multiple scenarios demonstrate the efficacy of our approach in utilizing credibility. Our model significantly outperforms various prevalent RAG approaches applied to both open and closed-source LLMs of diverse scales. Additionally, it exhibits robust resilience against noisy documents, maintaining high performance even as alternative strategies suffer sharp declines. All these results verify the effectiveness of the proposed CAG framework and corresponding training algorithm.
@@
Figure 2: Overview of data transformation framework. The training data is constructed by assigning credibility to contexts via multi-granularity credibility annotation (§3.1) and prompting LLM to produce credibility-guided explanations (§3.2). The processed data is used to instruction fine-tuning (§3.3) to endow the model with the ability for Credibility-aware Generation.
The main contributions of this study are summarized as follows  [^2] We uploaded the code and datasets as supplemental mate-rials, which will be openly released after accepting.:
@@
  • We propose a novel data transformation framework that transforms existing datasets into data annotated with credibility and guides models to generate responses based on credibility, thereby equipping the model with Credibility-aware Generation capability.
@@
  • Experimental evidences demonstrate that our model effectively understands and employs credibility to generate responses, significantly surpasses other RAG-based strategies, and maintains robustness despite the increasing noise in the context.
@@ 3.1 Multi-granularity Credibility Annotation
To cater to the varied requirements for credibility across different scenarios and enhance the model’s comprehension of credibility, we collect training data including open-domain QA, machine reading comprehension, and dialogue datasets and propose a multi-granularity credibility annotation method.
First, we divide the retrieved documents to create a multi-granularity corpus, encompassing sentence and document levels. Then, the retriever assesses the match between each retrieval unit and the query, assigning a relevance score, and classifies documents into three levels: high, medium, and low, using either equal count or equal interval methods. This approach of using levels instead of scores aims to simplify representation, thereby improving the model’s understanding and providing a certain degree of fault tolerance. Ultimately, we collect about 15k training data samples, all of which include documents with credibility annotations. The detailed composition of the training data is shown in the Appendix A.1.
@@ 3.2 Credibility-guided Explanation Generation
To facilitate the model’s comprehension and effective utilization of credibility, we employ LLMs to generate credibility-guided explanations for the answers.
Given the limitations of current LLMs in comprehending credibility effectively, we design chain-of-thought prompts to guide LLMs to generate credibility-guided explanations given questions, retrieved documents with credibility and golden answers. In this case, LLMs only need to generate coherent explanations based on the document containing the answers, without distinguishing between documents with different credibility to generate answers. The credibility-guided explanation obtained includes an analysis that integrates both the credibility and the content of the documents, rather than merely focusing on deriving the answer.
Considering the accessibility and advanced capabilities, we employ GPT-3.5 for the generation of explanations. In this way, we obtain high-quality answer explanations. Then, we replace the original answers in the training data with credibility-guided explanations to form a novel QA dataset. In this dataset, the inputs include questions and external documents annotated with credibility, while the outputs are credibility-guided explanations.
@@ 3.3 Instruction Fine-tuning
Through the two steps above, the training dataset obtained contains credibility, which can be used to facilitate arbitrary language models in gaining the capacity for CAG. We fine-tune the language model on this dataset to empower the model to discern and process information according to its credibility. As defined by Iyer et al. (2023), the loss function is as follows:
$\mathcal{L}(D_{\boldsymbol{x}};\theta)=-\sum_{i=1}^{N}\log p_{\theta}\left(\boldsymbol{y}_{i}\mid\left[\boldsymbol{x},\left\{[c_{i},d_{i}]\right\}_{i=1}^{|D_{\boldsymbol{x}}|}\right],\boldsymbol{y}_{<i}\right)$
@@ 4 Credibility-aware Generation Benchmark
$\bullet$ Open-domain QA aims to accurately answer questions on a wide variety of topics without being limited to any particular area. It encompasses a broad spectrum of real-world applications that urgently require the integration of external knowledge to enhance the LLMs’ ability to address queries. This scenario necessitates the ability to effectively identify and process noise information.
$\bullet$ Time-sensitive QA aims to give accurate and current answers. It poses a challenge for LLMs due to the dynamic internet information. The inevitable inclusion of outdated documents when incorporating external sources further complicates matters. Even with timestamps provided for documents, LLMs may erroneously prioritize outdated documents. This situation underscores the critical need for credibility in time-sensitive QA.
$\bullet$ Misinformation polluted QA aims to tackle the issue of ensuring accurate answers in an environment polluted with misinformation. It presents a substantial challenge to LLMs, attributed to the misuse of LLMs and the consequent proliferation of fake news and misinformation (Zhuo et al., 2023; Pan et al., 2023). Consequently, it is crucial to take into account the quality and credibility of any introduced external information. In the following, we will provide a detailed description of data construction for each scenario, and the statistics of CAGB are shown in the Table 1.
4.1 Credibility Assessment
We aim to establish a flexible credibility assessment mechanism that can be conveniently extended to consider additional factors and a broader range of application fields. In this benchmark, the credibility of the documents is evaluated by considering retrieval relevance, timeliness, and source reliability. Specifically, we establish a foundation based on retrieval relevance, then make adjustments according to timeliness, and finally integrate the reliability of the source to determine credibility. First, the retriever assigns relevance scores to documents based on query similarity. These relevance scores, which are distributed at equal intervals, enable to classify documents into three levels: high, medium, and low, collectively denoted as $R$. Subsequently, the temporal difference $T$ between the query time and document publication is calculated, downgrading $R$ if $T$ surpasses a threshold. The formula integrating relevance and timeliness is as follows:
$rt\_score(R,T)=\text{max}(R-\text{floor}(T/\text{threshold}),1)$
Following this, the source reliability, denoted $S$, is customized to specific scenarios, similarly divided into three levels. Finally, we combine these factors, adopting the lower level as the credibility and the formula is expressed as follows:
$Cred=\mathrm{min}(rt\_score(R,T),S)$
In this way, the document of high credibility are concurrently characterized by high relevance, timeliness and source reliability. More details about the assessment can be seen in the Appendix A.9.
@@ 4.2 Open-domain QA
Our research utilizes data from several challenging QA datasets with noisy documents. HotpotQA (Yang et al., 2018) and 2WikiMHQA (Ho et al., 2020) both require reasoning across multiple documents, and feature a high proportion of distracting documents. Importantly, the data we utilize from HotpotQA is extracted from the dev subset, whereas our training dataset is derived from the train subset. Musique (Trivedi et al., 2021) questions are of higher complexity, with up to 90% of distracting passages. ASQA (Stelmakh et al., 2022) is a long format QA dataset focused on ambiguous questions. RGB (Chen et al., 2023) is a specialized benchmark used for evaluating the capabilities of models in the RAG scenario, with noise robustness being one of its aspects. We assign credibility to the documents provided in the dataset in terms of retrieval relevance.
@@
Table 1: Statistics of CAGB, which includes 7 dataset derived from 3 scenarios.
@@ 4.3 Time-sensitive QA
In order to construct a diverse, high-quality, and up-to-date news dataset, we annotate 321 time-sensitive questions along with their corresponding dates. These questions originate from real-world scenarios, including news QA data from RealTime QA (Kasai et al., 2022), TAQA (Zhao et al., 2024), and questions adapted from news reports. To simulate the simultaneous occurrence of varied information on the Internet, we use Google search API to retrieve each query, selecting 3 relevant documents and 4 distracting documents. The distracting documents are either irrelevant to the query or outdated. This approach to document selection is crafted to emulate the intricate and heterogeneous nature of real-world information landscapes. Each news includes its publication date, thereby aiding in the evaluation of its timeliness. For document credibility annotation, we assess credibility based on relevance and time gap between the document’s publication and the posed question. We ensure the accuracy of the answers by manually annotating.
@@
ChatGPT 0.334 0.368 0.194 0.404 0.773 0.579 0.231
LLaMA-2-7B 0.280 0.312 0.160 0.268 0.753 0.433 0.179
Vicuna-7B 0.278 0.296 0.116 0.358 0.677 0.567 0.229
Mistral-7B-Instruct 0.288 0.270 0.106 0.300 0.713 0.598 0.204
LLaMA-2-13B 0.366 0.370 0.164 0.321 0.820 0.495 0.204
LLaMA-2-70B 0.418 0.390 0.256 0.316 0.823 0.526 0.430
vanilla IFT 0.324 0.245 0.270 0.157 0.650 0.592 0.329
@@
ChatGPT 0.396 0.394 0.216 0.388 0.790 0.632 0.427
LLaMA-2-7B 0.302 0.376 0.200 0.375 0.730 0.526 0.265
Vicuna-7B 0.355 0.306 0.164 0.494 0.757 0.620 0.275
Mistral-7B-Instruct 0.338 0.334 0.166 0.414 0.790 0.741 0.373
LLaMA-2-13B 0.370 0.372 0.180 0.390 0.823 0.561 0.308
LLaMA-2-70B 0.422 0.504 0.320 0.388 0.833 0.570 0.306
vanilla IFT 0.348 0.448 0.276 0.304 0.663 0.720 0.344
@@
ChatGPT 0.422 0.402 0.182 0.440 0.807 0.673 0.408
LLaMA-2-7B 0.376 0.176 0.140 0.394 0.713 0.486 0.213
Vicuna-7B 0.349 0.266 0.091 0.490 0.740 0.642 0.279
Mistral-7B-Instruct 0.274 0.268 0.102 0.463 0.797 0.679 0.315
LLaMA-2-13B 0.360 0.384 0.164 0.385 0.803 0.520 0.227
@@
vanilla IFT 0.372 0.334 0.204 0.305 0.663 0.589 0.383
CAG-7B (ours) 0.509 0.578 0.340 0.496 0.897 0.826 0.442
CAG-13B (ours) 0.514 0.604 0.408 0.525 0.917 0.829 0.483
CAG-mistral-7B (ours) 0.502 0.540 0.384 0.505 0.900 0.835 0.613
Table 2: Model performance in our CAGB benchmark. The best/second best scores in each dataset are bolded/underlined. Our models substantially outperform previous strategies across all 3 scenarios in CAGB. The results shown for EvolvingTempQA and RGB are at noise_ratio setting of 0.8, while NewsPollutedQA is at noise_ratio setting of 0.75. The results of other metrics on the ASQA dataset are shown in the Appendix A.6.
4.4 Misinformation Polluted QA
To simulate the generation of fake news, we first generated a claim using LLMs, based on a question and a randomly selected incorrect option. This process transforms the question and incorrect option into a deceptive statement. Subsequently, we choose GPT-3.5 and Qwen (Bai et al., 2023) as the generators for fake news, guiding them to generate texts of varying styles based on the claim, including news style and Twitter style. The prompts used and examples are detailed in the Appendix A.15. The fictitious news articles produced by LLMs, due to their authenticity being deliberately compromised, are classified as having low source reliability. Conversely, news articles from reputable news websites are considered to possess high source reliability. We set the ratio of fake news at 0.5, 0.67, and 0.75 to evaluate the robustness of model against misinformation under various levels of pollution.
By simulating fake news generation, we create a misinformation polluted QA dataset in the news domain, named NewsPollutedQA.